# Beyond Thumbs Up/Down: Untangling Challenges of Fine-Grained Feedback for Text-to-Image Generation

### v1

@@ 8.1. Which Attributes to Elicit?
Additionally, it is not enough just to have the “right” attributes. In practice, elicitation needs to balance informativity and cost. We already see that attributes may take different amounts of time to annotate in Experiment 1 (see see Table 2) and in Experiment 2, we demonstrate that, in simulation, adding attributes is not always valuable. Interdisciplinary works that straddle AI, cognitive science, and human-computer interaction are already exploring the impact of requiring humans to provide feedback on many attributes, noting that such a practice can overwhelm cognitive load and risk bringing more error into downstream modeling (Sucholutsky et al., 2023; Ramaswamy et al., 2022; Barker et al., 2023). Indeed, we do not want to waste annotations on attributes where users are highly unsure (though future work can explore the benefits from learning with uncertainty at feedback time (Collins et al., 2023a)). And further, we already see in our human studies that annotators spend substantially different time annotating some attributes over others. Nicely, the CBM model class naturally supports the implementation and study of cost-aware acquisition strategies for human feedback  (Chauhan et al., 2022; Sheth et al., 2022; Shin et al., 2023; Espinosa Zarlenga et al., 2024). We see promise in the hybridization of elicitation development for such models and the determination of which attributes to elicit.
8.2. Reward Model Structure
@@
Yet, perhaps CBMs are not the best model structure for fine-grained feedback. Indeed, recent work has raised questions about the ability of CBM-based systems to effectively handle rich, soft-labeled feedback if not explicitly trained to do so Collins et al.. This highlights the importance of considering the interplay between feedback type and model architecture when designing systems for human-in-the-loop adaptation. We reiterate that our work is a preliminary exploration of ways to learn reward models from fine-grained feedback. It is likely that alternate modeling choices induce different cost-benefit analyses on the value of learning from fine-grained feedback; we look forward to future works that explore such possibilities. For instance, several other approaches have been proposed e.g.  (Rame et al., 2024; Liang et al., 2023). We see the design of model architectures which incorporate information efficiently from granular feedback, and can flexibly grow to handle new dimensions (e.g, if we learn that a new attribute actually matters more to annotators that we had not previously modeled), as ripe for future work. Moreover, the image and text embeddings we considered in our work were always fixed. It is possible that different choices of embedding, or even jointly learning embeddings, may improve performance and perhaps salvage the utility of a CBM-based architecture.
8.3. Accessible, Efficient Evaluation
@@
8.4. Human vs. Model Feedback
Computational overhead is not the only challenge: we are also limited by the elicitation of feedback itself. Eliciting information from humans can be expensive. Here, our granular feedback was derived from an AI system, not humans. A natural question is how well our simulated feedback here actually correlates to human judgments. It is possible that our null results in Experiment 1 stem from a mismatch between human and model judgments over the granular attributes, either or both along the image quality and text-image alignment dimensions. While there is a push to employ AI-generated feedback rather than humans for scalable generative evaluation  (Wu et al., 2023b; Gilardi et al., 2023), it is essential to understand where such feedback may diverges from human expectations.
9. Conclusion
@@ Acknowledgments
Appendix A Appendix
A.1. Image and Text Embeddings
Image and text inputs are represented as dense embedding vectors. We use CLIP to extract embeddings for text captions (Radford et al., 2021). Through preliminary experimentation, we found the frozen CLIP embeddings has been shown to poorly capture aesthetic properties of images; as such, we opted to extract the output of the first layer of the LAOIN Aeshetics model (Schuhmann et al., 2022) as our image embeddings. Embeddings are concatenated across modalities for Experiment 1; only the image embeddings are fed as used as input for training and inference in Experiment 2. Image and text embeddings are frozen for all models. Future work could explore the impact of jointly fine-tuning the embeddings and predicting granular feedback.
A.2. Reward Models
As introduced in Section 3, we run a two-phased training procedure for fine-grained reward models. We first train a mapping from the input embeddings (described above) to individual attributes; this takes the form of a multi-headed MLP. We then learn a simple linear aggregator over the outputs of the multi-headed MLP. All stages leverage binary classifiers for a form of feedback (specifically, multi-class binary classifiers when we have multiple attributes); the input to the Stage 2 linear aggregator for all settings is the sigmoided logit from Stage 1. The coarse-grained baseline only involves stage one (we directly map from the input embeddings to the coarse score); i.e., the coarse-grained model is not a CBM.
@@ A.2. Reward Models
A.3. Additional Details on Forms of Fine-Grained Feedback
A.3.1. Image Quality
@@
A.3.2. Prompt Alignment
We build on the $VQ^{2}$ method developed in (Yarom et al., 2023) to measure the alignment between a prompt and the image. As discussed, $VQ^{2}$ takes as input an image and a prompt (e.g., “green dog to the left of the river”) and generates a series of binary questions that the image ought to address if fit for the prompt (e.g., “is there a dog?”, “is the dog green?”, “is the dog running along the left of the river?”). Each question is then assessed against the image, wherein the probability that the question can be answered as “Yes” is computed. The mean of the “Yes” probabilities forms the final $VQ^{2}$ score.
Here, we construct four scores to reflect different ways in which an image may be aligned to a prompt. An image may be aligned in its: 1) representations of objects / nouns (e.g., “is there a dog?”), 2) attributes / adjectives (“is the dog green?”), 3) actions / verbs (“is the dog running?”), and 4) relations (“is the dog to the left of the river?”). We categorize each of the questions generated by $VQ^{2}$ into one of these categories using a custom semantic parser built from spaCy  (Honnibal et al., 2020) with hand-crafted rules to catch exceptions. Two authors from our author team manually inspected hundreds of the categorizations to affirm their quality – while the parsing was generally sensible – we note that it is not perfect and likely could be improved in future work. We then average the $VQ^{2}$ scores for all questions grouped in a category, which are then thresholded into a binary aligned/not aligned which we use as feedback. Image-prompt pairs for which $VQ^{2}$ does not generate a question for a particular class are binarized into the positive (i.e., aligned) class as we care more about cases which are mis-aligned along an attribute. We encourage future work to improve both the question generation, classification, and answer categorization.
A.4. Additional Details on Data Processing
We form preference judgements by aggregating over the contextually-annotated images from  (Dutta et al., 2024). We apply simple averaging, where each annotation is weighted equally – alternate weighting schemes could be worth exploring in the future, as well as a breakdown by the context. We split the data along the prompts, as there are four different images per prompt, each annotated with human scores.
@@ A.4. Additional Details on Data Processing
A.5. Additional Details on Rejection Sampling
A.6. Stimuli Generation
@@
A.7. Reward Model Scoring and Selection
@@
A.8. Human Study: SxS Evals
@@
A.9. Visualizing the Learned Aggregators
@@
Figure 6. Attribute weights learned for the prompt-aware setting; most weight is placed on attributes scoring prompt-image alignment.
Figure 7. Attribute weights learned for the decision tree setting. The aggregator appropriately learns the importance and direction of the attributes which make up the leaves of the tree.
Figure 8. Attribute weights learned for the decision tree setting. The aggregator learns to focus on attributes semantically related to those which form the decision tree (e.g., captivating versus visual compelling).
A.10. Samples Scored by Fine-Grained Models
@@
Figure 9. Images scored by the trained fine-grained reward model. Images in the top row are those which are rated high on an attribute (noted on the y axis), and in the bottom row, rated low by the reward model in terms of that attribute. Note, “compelling” here is “visually compelling”.

### v2

@@ Results
Additionally, it is not enough just to have the “right” attributes. In practice, elicitation needs to balance informativity and cost. We already see that attributes may take different amounts of time to annotate in Experiment 1 (see Table 2) and in Experiment 2, we demonstrate that, in simulation, adding attributes is not always valuable. Interdisciplinary works that straddle AI, cognitive science, and human-computer interaction are already exploring the impact of requiring humans to provide feedback on many attributes, noting that such a practice can overwhelm cognitive load and risk bringing more error into downstream modeling (Sucholutsky et al. 2023; Ramaswamy et al. 2022; Barker et al. 2023). Indeed, we do not want to waste annotations on attributes where users are highly unsure (though future work can explore the benefits from learning with uncertainty at feedback time (Collins et al. 2023a)). And further, we already see in our human studies that annotators spend substantially different time annotating some attributes over others. Nicely, the CBM model class naturally supports the implementation and study of cost-aware acquisition strategies for human feedback  (Chauhan et al. 2022; Sheth et al. 2022; Shin et al. 2023; Espinosa Zarlenga et al. 2024). We see promise in the hybridization of elicitation development for such models and the determination of which attributes to elicit.
Reward Model Structure
@@
Yet, perhaps CBMs are not the best model structure for fine-grained feedback. Indeed, recent work has raised questions about the ability of CBM-based systems to effectively handle rich, soft-labeled feedback if not explicitly trained to do so Collins et al.. This highlights the importance of considering the interplay between feedback type and model architecture when designing systems for human-in-the-loop adaptation. We reiterate that our work is a preliminary exploration of ways to learn reward models from fine-grained feedback. It is likely that alternate modeling choices induce different cost-benefit analyses on the value of learning from fine-grained feedback; we look forward to future works that explore such possibilities. For instance, several other approaches have been proposed e.g.  (Rame et al. 2024; Liang et al. 2023). We see the design of model architectures which incorporate information efficiently from granular feedback, and can flexibly grow to handle new dimensions (e.g, if we learn that a new attribute actually matters more to annotators that we had not previously modeled), as ripe for future work. Moreover, the image and text embeddings we considered in our work were always fixed. It is possible that different choices of embedding, or even jointly learning embeddings, may improve performance and perhaps salvage the utility of a CBM-based architecture.
Accessible, Efficient Evaluation
@@
Human vs. Model Feedback
Computational overhead is not the only challenge: we are also limited by the elicitation of feedback itself. Eliciting information from humans can be expensive. Here, our granular feedback was derived from an AI system, not humans. A natural question is how well our simulated feedback here actually correlates to human judgments. It is possible that our null results in Experiment 1 stem from a mismatch between human and model judgments over the granular attributes, either or both along the image quality and text-image alignment dimensions. While there is a push to employ AI-generated feedback rather than humans for scalable generative evaluation  (Wu et al. 2023a; Gilardi, Alizadeh, and Kubli 2023), it is essential to understand where such feedback may diverges from human expectations (Collins et al. 2024).
Conclusion
@@ Acknowledgements
Appendix
Image and Text Embeddings
Image and text inputs are represented as dense embedding vectors. We use CLIP to extract embeddings for text captions (Radford et al. 2021). Through preliminary experimentation, we found the frozen CLIP embeddings has been shown to poorly capture aesthetic properties of images; as such, we opted to extract the output of the first layer of the LAOIN Aeshetics model (Schuhmann et al. 2022) as our image embeddings. Embeddings are concatenated across modalities for Experiment 1; only the image embeddings are fed as used as input for training and inference in Experiment 2. Image and text embeddings are frozen for all models. Future work could explore the impact of jointly fine-tuning the embeddings and predicting granular feedback.
Reward Models
As introduced in Section Problem Setting, we run a two-phased training procedure for fine-grained reward models. We first train a mapping from the input embeddings (described above) to individual attributes; this takes the form of a multi-headed MLP. We then learn a simple linear aggregator over the outputs of the multi-headed MLP. All stages leverage binary classifiers for a form of feedback (specifically, multi-class binary classifiers when we have multiple attributes); the input to the Stage 2 linear aggregator for all settings is the sigmoided logit from Stage 1. The coarse-grained baseline only involves stage one (we directly map from the input embeddings to the coarse score); i.e., the coarse-grained model is not a CBM.
@@
Additional Details on Forms of Fine-Grained Feedback
Image Quality
@@
Figure 5: Estimated similarity between PaLI scores for different attributes. We depict the proportion of images in the train set for which PaLI marks an image as having the same attribute (e.g., the cell blurry and malformed highlights that PaLI marks an image as blurry and malformed, or not blurry and not malformed for 71% of the examples). Darker red means higher level of similarity in scores; yellow represents lower similarity.
Prompt Alignment
We build on the $VQ^{2}$ method developed in (Yarom et al. 2023) to measure the alignment between a prompt and the image. As discussed, $VQ^{2}$ takes as input an image and a prompt (e.g., “green dog to the left of the river”) and generates a series of binary questions that the image ought to address if fit for the prompt (e.g., “is there a dog?”, “is the dog green?”, “is the dog running along the left of the river?”). Each question is then assessed against the image, wherein the probability that the question can be answered as “Yes” is computed. The mean of the “Yes” probabilities forms the final $VQ^{2}$ score.
Here, we construct four scores to reflect different ways in which an image may be aligned to a prompt. An image may be aligned in its: 1) representations of objects / nouns (e.g., “is there a dog?”), 2) attributes / adjectives (“is the dog green?”), 3) actions / verbs (“is the dog running?”), and 4) relations (“is the dog to the left of the river?”). We categorize each of the questions generated by $VQ^{2}$ into one of these categories using a custom semantic parser built from spaCy  (Honnibal et al. 2020) with hand-crafted rules to catch exceptions. Two authors from our author team manually inspected hundreds of the categorizations to affirm their quality – while the parsing was generally sensible – we note that it is not perfect and likely could be improved in future work. We then average the $VQ^{2}$ scores for all questions grouped in a category, which are then thresholded into a binary aligned/not aligned which we use as feedback. Image-prompt pairs for which $VQ^{2}$ does not generate a question for a particular class are binarized into the positive (i.e., aligned) class as we care more about cases which are mis-aligned along an attribute. We encourage future work to improve both the question generation, classification, and answer categorization.
Additional Details on Data Processing
We form preference judgements by aggregating over the contextually-annotated images from  (Dutta et al. 2024). We apply simple averaging, where each annotation is weighted equally – alternate weighting schemes could be worth exploring in the future, as well as a breakdown by the context. We split the data along the prompts, as there are four different images per prompt, each annotated with human scores.
@@
Additional Details on Rejection Sampling
Stimuli Generation
@@
Reward Model Scoring and Selection
@@
Human Study: SxS Evals
@@
Visualizing the Learned Aggregators
@@
Figure 6: Attribute weights learned for the prompt-aware setting; most weight is placed on attributes scoring prompt-image alignment.
Figure 7: Attribute weights learned for the decision tree setting. The aggregator appropriately learns the importance and direction of the attributes which make up the leaves of the tree.
Figure 8: Attribute weights learned for the decision tree setting. The aggregator learns to focus on attributes semantically related to those which form the decision tree (e.g., captivating versus visual compelling).
Samples Scored by Fine-Grained Models
@@
Figure 9: Images scored by the trained fine-grained reward model. Images in the top row are those which are rated high on an attribute (noted on the y axis), and in the bottom row, rated low by the reward model in terms of that attribute. Note, “compelling” here is “visually compelling”.